# Systematic search for cool hypervelocity stars in multi-epoch WISE data

### v1

@@ 4 Conclusions
We performed a systematic search for the objects with high proper motions in W2 band WISE data using recently released unTimely epochal catalogue, and successfully identified 3245 objects moving faster than 0.6′′/year. Our estimations of their proper motions are in a good agreement with Gaia DR3 measurements where they are available. We did not detect any object moving faster than approx. 10′′/year, but were able to successfully recover both the fastest-moving known brown dwarf WISEA J085510.74-071442.5 and most of other known rapidly moving normal stars and nearby cool dwarfs. We also identified 42 objects that do not have associated records in Simbad database, 32 of them not previously published. We assembled multi-wavelength information for them by cross-matching their trajectories with large optical and infrared catalogues, and used their colors to estimate their spectral types, distances and projected velocities. At least 15 of them are reliable brown dwarf candidates with estimated distances closer than 50 pc and types later than T0.
@@ Acknowledgements.
This work was co-funded by the EU and supported by the Czech Ministry of Education, Youth and Sports through the project CZ.02.01.01/00/22_008/0004596. This research was partially funded by the Ministry of Science and Higher Education of the Russian Federation, according to the research project 13.2251.21.0177 (075-15-2022-1228). This work has benefitted from The UltracoolSheet at http://bit.ly/UltracoolSheet, maintained by Will Best, Trent Dupuy, Michael Liu, Aniket Sanghi, Rob Siverd, and Zhoujian Zhang, and developed from compilations by Dupuy & Liu (2012, ApJS, 201, 19), Dupuy & Kraus (2013, Science, 341, 1492), Deacon et al. (2014, ApJ, 792, 119), Liu et al. (2016, ApJ, 833, 96), Best et al. (2018, ApJS, 234, 1), Best et al. (2021, AJ, 161, 42), Sanghi et al. (2023, ApJ, 959, 63), and Schneider et al. (2023, AJ, 166, 103). This publication makes use of VOSA, developed under the Spanish Virtual Observatory (https://svo.cab.inta-csic.es) project funded by MCIN/AEI/10.13039/501100011033/ through grant PID2020-112949GB-I00. VOSA has been partially updated by using funding from the European Union’s Horizon 2020 Research and Innovation Programme, under Grant Agreement No 776403 (EXOPLANETS-A)References
- Abbott et al. (2021) Abbott, T. M. C., Adamów, M., Aguena, M., et al. 2021, ApJS, 255, 20
@@
- Burgasser et al. (2024) Burgasser, A. J., Gerasimov, R., Kremer, K., et al. 2024, arXiv e-prints, arXiv:2407.08578
@@
- Geballe et al. (2002) Geballe, T. R., Knapp, G. R., Leggett, S. K., et al. 2002, ApJ, 564, 466
@@
- Lodieu et al. (2014) Lodieu, N., Pérez-Garrido, A., Béjar, V. J. S., et al. 2014, Astronomy and Astrophysics, 569, A120
@@
- Tinney et al. (2018) Tinney, C. G., Kirkpatrick, J. D., Faherty, J. K., et al. 2018, ApJS, 236, 28

### v2

@@ 4 Conclusions
We performed a systematic search for the objects with high proper motions in band WISE data using recently released unTimely epochal catalogue, and successfully identified 21,885 objects moving faster than 0.3′′/year. Our estimations of their proper motions are in a good agreement with Gaia DR3 measurements where they are available. We did not detect any object moving faster than approx. 10′′/year, but were able to successfully recover both WISEA J085510.74-071442.5 brown dwarf with largest known proper motion and most of other known rapidly moving normal stars and nearby ultracool dwarfs. We also identified 258 objects that do not have associated records in SIMBAD database or Gaia catalogues, and had not been published previously. We assembled multi-wavelength information for them by cross-matching their trajectories with large optical and infrared catalogues, and used their colors to estimate their effective temperatures, spectral types, distances and projected velocities. 252 are compatible with being ultracool dwarfs, and at least 33 of them are reliable T brown dwarf candidates with estimated distances closer than 40 pc and effective temperatures colder than 1300 K.
@@ Acknowledgements.
This work was co-funded by the EU and supported by the Czech Ministry of Education, Youth and Sports through the project CZ.02.01.01/00/22_008/0004596. This work has benefitted from The UltracoolSheet at http://bit.ly/UltracoolSheet, maintained by Will Best, Trent Dupuy, Michael Liu, Aniket Sanghi, Rob Siverd, and Zhoujian Zhang, and developed from compilations by Dupuy & Liu (2012, ApJS, 201, 19), Dupuy & Kraus (2013, Science, 341, 1492), Deacon et al. (2014, ApJ, 792, 119), Liu et al. (2016, ApJ, 833, 96), Best et al. (2018, ApJS, 234, 1), Best et al. (2021, AJ, 161, 42), Sanghi et al. (2023, ApJ, 959, 63), and Schneider et al. (2023, AJ, 166, 103). This publication makes use of VOSA, developed under the Spanish Virtual Observatory (https://svo.cab.inta-csic.es) project funded by MCIN/AEI/10.13039/501100011033/ through grant PID2020-112949GB-I00. VOSA has been partially updated by using funding from the European Union’s Horizon 2020 Research and Innovation Programme, under Grant Agreement No 776403 (EXOPLANETS-A)References
@@
- Allard (2014) Allard, F. 2014, in IAU Symposium, Vol. 299, Exploring the Formation and Evolution of Planetary Systems, ed. M. Booth, B. C. Matthews, & J. R. Graham, 271–272
- Andrae et al. (2023) Andrae, R., Fouesneau, M., Sordo, R., et al. 2023, A&A, 674, A27
@@
- Brown (2015) Brown, W. R. 2015, ARA&A, 53, 15
- Burgasser et al. (2024a) Burgasser, A. J., Bezanson, R., Labbe, I., et al. 2024a, ApJ, 962, 177
- Burgasser et al. (2024b) Burgasser, A. J., Schneider, A. C., Meisner, A. M., et al. 2024b, arXiv e-prints, arXiv:2411.01378
@@
- Burningham et al. (2017) Burningham, B., Marley, M. S., Line, M. R., et al. 2017, Monthly Notices of the Royal Astronomical Society, 470, 1177
@@
- Charnay et al. (2018) Charnay, B., Bézard, B., Baudino, J. L., et al. 2018, The Astrophysical Journal, 854, 172
@@
- Gerasimov et al. (2024) Gerasimov, R., Bedin, L. R., Burgasser, A. J., et al. 2024, ApJ, 971, 65
- Gizis et al. (2022) Gizis, J. E., Yoachim, P., Jones, R. L., Hilligoss, D., & Ji, J. 2022, ApJS, 263, 23
- Hainline et al. (2024) Hainline, K. N., Helton, J. M., Johnson, B. D., et al. 2024, ApJ, 964, 66
@@
- Hsu et al. (2024) Hsu, C.-C., Burgasser, A. J., Theissen, C. A., et al. 2024, ApJS, 274, 40
- Humphreys et al. (2020) Humphreys, A., Schapera, N., Meisner, A. M., et al. 2020, in Astronomical Society of the Pacific Conference Series, Vol. 525, Compendium of Undergraduate Research in Astronomy and Space Science, ed. J. B. Jensen, J. Barnes, & B. Wardell, 57
- Khandrika et al. (2013) Khandrika, H. G., Burgasser, A. J., Melis, C., et al. 2013, The Astronomical Journal, 145
@@
- Kota et al. (2022) Kota, T., Kirkpatrick, J. D., Caselden, D., et al. 2022, AJ, 163, 116
- Kroupa et al. (2024) Kroupa, P., Gjergo, E., Jerabkova, T., & Yan, Z. 2024, arXiv e-prints, arXiv:2410.07311
- Kroupa et al. (2013) Kroupa, P., Weidner, C., Pflamm-Altenburg, J., et al. 2013, in Planets, Stars and Stellar Systems. Volume 5: Galactic Structure and Stellar Populations, ed. T. D. Oswalt & G. Gilmore, Vol. 5, 115
@@
- Langeveld et al. (2024) Langeveld, A. B., Scholz, A., Mužić, K., et al. 2024, AJ, 168, 179
@@
- Luhman (2024) Luhman, K. L. 2024, AJ, 168, 159
- Malkov (2023) Malkov, O. Y. 2023, Astronomy Reports, 67, 288
- Marley & Robinson (2015) Marley, M. S. & Robinson, T. D. 2015, Annual Review of Astronomy and Astrophysics, 53, 279
@@
- Tan & Showman (2019) Tan, X. & Showman, A. P. 2019, The Astrophysical Journal, 874, 111
- Tremblin et al. (2019) Tremblin, P., Padioleau, T., Phillips, M. W., et al. 2019, The Astrophysical Journal, 876, 144
- Wang et al. (2022) Wang, Y.-F., Luo, A. L., Chen, W.-P., et al. 2022, A&A, 660, A38